Solve using the quadratic formula:
x^2-7x-1=-7

Answer by checkley71(8403)

X^2-7X-1=-7
X^2-7X-1+7=0
X^2-7X+6=0
(X-6)(X-1)=0
X-6=0
X=6 ANSWER.
X-1=0
X=1 ANSWER.